OpenStudy (anonymous):

logarithm question:

OpenStudy (anonymous):

ill post it:

OpenStudy (anonymous):

OK... keep us updated...

OpenStudy (anonymous):

\[5^\log x^5 =x^9\]

OpenStudy (anonymous):

evaluate and fins the exact answer for x

OpenStudy (anonymous):

I'm not quite sure how to solve this one either but we can get it down to this: \[5^{\ln x}x^5 = x^9 simplifies \to 5^{\ln x} = x^4\] Then if we take the log of each side: \[\ln 5^{\ln x } =\ln x^4 simplifies \to \ln x * \ln 5 = 4 \ln x\] And as we see ln 5 will never equal 4 then we must get ln x to equal 0 and then we will have 0 = 0. So ln x = 0 means x = 1. Anyone have a more exact approach to this?

OpenStudy (anonymous):

thanks, but my teacher showed me me the correct approach :) (the answer was x= 5^1/3 btw)